Calculate the minimum deviation produced by a 60° glass prism if the refractive index of the glass is

    1.50. What is the angle of incidence at which minimum deviation occurs?


1
Expert's answer
2021-03-08T08:43:42-0500

Answer

Minimum angle of reflection should be half of prism angle

"r=\frac{A}{2}=30°"

"n_1sini=n_2sinr"

So

"i=sin^{-1}(\frac{n_2sinr}{n_1}) \\i=sin^{-1}(\frac{(1.5)sin30°}{1})=48.6°"